Trying to help my daughter with her math...really rusty could use some help so I can help her....here is the problem the projection screen in the school auditorium is 5 times as long and 5 times as wide as the screen in the library.. the screen in the library is 4 feet long with a perimeter of 14 feet. What is the perimeter of the screen in the auditorium? Any help would greatly appreciated :)

Answer by Alan3354(69443)

the projection screen in the school auditorium is 5 times as long and 5 times as wide as the screen in the library.. the screen in the library is 4 feet long with a perimeter of 14 feet. What is the perimeter of the screen in the auditorium?
---------------
Perimeter is a linear function.
5 times the dimensions --> 5x the perimeter --> 70 feet
----------
Or,
the screen in the library is 4 feet long with a perimeter of 14 feet.
P = 2W + 2L = 14
L + W = 7
4 + W = 7
W = 3
-----
Perimeter in the auditorium is
2*5*L + 2*5*W
= 10L + 10W = 10*4 + 10*3
= 70 feet